OpenStudy (anonymous):

A car with a velocity of 30 m/s is accelerated uniformly at the rate of 2.0 m/s2 for 10 s. What is its final velocity?

OpenStudy (xishem):

\[v=v_0+at\]

OpenStudy (anonymous):

the first thing you want to do is determine the given Vi=30m/s A=2.0m/s2 T=10s after that set up an equation, since you want to find the final velocity you can use the Vf=Vi+at plug in all the given number Vf=30m/s+(2.0 m/s2)(10s) so the Vf is equal to 50m/s hope that's helpful.